A solution of saltwater is 10% salt and weighs 50 lbs. How much salt is added to produce a solution that is 25% salt (in lbs)?
Answer by checkley71(8403)

.1*50+X=.25(50+X)
5+X=12.5+.25X
X-.25X=12.5-5
.75X=7.5
X=7.5/.75
X=10 POUNDS OF SALT ARE NEEDED.
PROOF.
.1*50+10=.25(50+10)
5+10=.25(60)
15=15